OpenStudy (anonymous):

x^(2/3) - 100=0 Help me solve this please!

OpenStudy (anonymous):

\[x^{2/3}-100=0\] \[x^{2/3}=100\] \[(x^{2/3})^3=100^3\] \[x^2=1,000,000\] \[\sqrt{x^2}=\sqrt{1,000,000}\] \[x=1000\]

OpenStudy (anonymous):

Thank you! that's what i thought... i put x=1000 as an answer on my test and i got marked wrong, as well as many of my friends with the same answer.

OpenStudy (anonymous):

I think you got it marked wrong, which I left this out of the answer by mistake, but when you are solving for x, after you take the square root, x=+/- 1000 because 1000 and -1000 both satisfy the equation. Sorry bout that.

OpenStudy (anonymous):

ohh that makes sense. whoops, oh well. thanks for your help though